The children at a local daycare planted a rectangular vegetable garden with a perimeter of 39 m. The length is twice the width. Find the dimensions of the garden.

Found 2 solutions by ikleyn, mathsolverplus:
Answer by ikleyn(52782)

.
Perimeter = W + 2W + W + 2W = 39 meters.

6W = 39 meters.

Solve for W, which is the width.

Then evaluate 2W, which is the length.


Answer by mathsolverplus(88)

Length = 2 (width)
Perimeter of a rectangle = 2(length + width) = 39
Let width = w and substitute length into the equation above ^:
2*(2w+w))=39
2*(3w) = 39
6w = 39
w = 39/6 = 13/2m
length = 2w = 2(13/2) = 13m
Dimensions of the garden:
length = 13m
width = 6.5m
If you need additional help, please message us through social media!!! We can't reply your comments here unfortunately :(
Instagram (@mathsolver.plus) : https://www.instagram.com/mathsolver.plus/
Twitter (@MathSolverPlus): https://twitter.com/MathSolverPlus
Snapchat (@mathsolverplus)
Email: mathsolverplus@gmail.com